OpenStudy (anonymous):

The formula T=11.0/d+0.320 gives the number of threads per inch in a standard machine screw which is d inches in diameter. Find the approximate value of T when d= .178 inch???

OpenStudy (opcode):

The formula \(T=\dfrac{11.0}{d+0.320}\) gives the number of threads per inch in a standard machine screw which is d inches in diameter. Find the approximate value of T when d= .178 inch. Step 1: Input known values: \(T=\dfrac{11.0}{.178+0.320}\) Step 2: Simplify: \(T=\dfrac{11.0}{0.498}\) Step 3: Simplify even more: \(T = 22.0884\) Here is a neat tip: Try crossing out unneeded information in word problems: The formula \(T=\dfrac{11.0}{d+0.320}\) \(\cancel{\text{gives the number of threads per inch in a standard machine screw which is d inches in diameter.}}\) Find the approximate value of T when d= .178 inch.